# Entropy-driven entanglement forging

### v1

@@ II.3 Nuclear shell model
In addition, nuclear entanglement can also be partly understood in terms of subshell closures and occupation numbers. Nuclei with proton or neutron orbitals mostly empty or fully-occupied present very low entanglement of these orbitals with the rest. This feature motivates to apply entanglement forging in an iterative fashion: firstly, using a proton-neutron partition; and secondly using an energy-based subpartition, separating the higher- and lower-energy single-particle states for each proton and neutron subsystem. Figure 4 illustrates the two partitions and the corresponding cuts used for 28Ne and 60Ti. For the latter nucleus, a bipartition into subsystems of the same size requires separating the energy-degenerate single-particle states of the $1p_{3/2}$ orbital.
@@ III Entropy-driven entanglement forging with ADAPT-VQE
ADAPT-VQE [54, 55, 56, 57, 58] is a variational quantum algorithm which updates iteratively a user-defined ansatz, rather than optimizing a fixed number of parameters as the UCC-VQE [20, 19, 3]. Each iteration $k$ adds a new unitary operator, $A_{k}=e^{i\theta_{k}T_{k}}$, to the ansatz, with a new parameter, $\theta_{k}$, and a hermitian operator from a predefined pool, $T_{k}$. Then, all parameters, $\bm{\theta}=\{\theta_{1},\cdots,\theta_{k}\}$, are optimized simultaneously to minimize the energy,
@@
This optimization starts with the values for $\theta_{1},\cdots,\theta_{k-1}$ from the minimization of the previous $k-1$ iteration and $\theta_{k}=0$. The operator $T_{k}$ is chosen such that the energy gradient with respect to $\theta_{k}$,
@@
is maximum. Thus, by starting from the minimum in the $(k-1)$-dimensional parameter space and maximizing the gradient in the new dimension, the algorithm gradually increases the parameter space while always decreasing the energy. ADAPT-VQE was originally proposed to solve the electronic structure of the ground state of molecules [59, 60], and since then it has been applied to a broad variety of quantum many-body systems, including lattice quantum chromodynamics [61, 62], the NSM [18, 25], algebraic nuclear models [18, 63] and nuclear pairing [64]. In addition to the novel FH simulations, we build on our previous work [25, 31] and optimize the ADAPT-VQE algorithm to solve NSM ground states with fewer quantum resources by using EDEF.
@@
with the restriction that they conserve the Hamiltonian symmetries: the spin for the FH and $M,T_{z}$ for the NSM. For the FH, in addition, we also include one-body operators,
$T_{s}^{r}=i(a_{r}^{{\dagger}}a_{s}-a_{s}^{{\dagger}}a_{r}).$ (12)
@@
Figure 6: Infidelity as a number of ADAPT-VQE iterations for the same FH model parameters (middle hopping $t_{m}$, interaction $U$) as in Fig. 3. ADAPT-VQE without entanglement forging (solid red lines) is compared with one layer of EDEF (labelled “1 cut”, dashed blue lines). Horizontal lines indicate the infidelity minimum set by the Schmidt-decomposition cutoff.
@@
As a reference state, with the standard ADAPT-VQE we choose the lowest-energy Slater determinant in the Fock basis of the particular FH lattice or nucleus, as it can be implemented with only one-qubit gates with the Jordan-Wigner mapping. With EDEF, however, we can exploit other physical arguments, for example the degeneracies in the Schmidt decomposition. In the case of the FH model, the best choice is a symmetric distribution of up and down spins between the left ($l$) and right ($r$) parts of the lattice:
$|\psi\rangle=$ $\lambda_{0}|l_{\uparrow\downarrow}\rangle\otimes|r_{\uparrow\downarrow}\rangle+\lambda_{1}\bigg{(}|l_{\uparrow\uparrow\downarrow}\rangle\otimes|r_{\downarrow}\rangle+|l_{\downarrow}\rangle\otimes|r_{\uparrow\uparrow\downarrow}\rangle$ (13)
$+|l_{\uparrow\downarrow\downarrow}\rangle\otimes|r_{\uparrow}\rangle+|l_{\uparrow}\rangle\otimes|r_{\uparrow\downarrow\downarrow}\rangle\bigg{)}.$
@@
In turn, for the NSM we use eigenstates of the $J_{z}$ operator in the proton ($p_{M}$) and neutron ($n_{M}$) sectors, with the pattern described in Sec. II.3,
$|\psi\rangle=$ $\lambda_{0}|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle+\lambda_{1}\bigg{(}|p_{-2}\rangle\otimes|n_{2}\rangle+|p_{-1}\rangle\otimes|n_{1}\rangle$ (14)
$+|p_{0}\rangle\otimes|n_{0}\rangle+|p_{1}\rangle\otimes|n_{-1}\rangle+|p_{2}\rangle\otimes|n_{-2}\rangle\bigg{)}.$
@@
The operator pool preserves the quantum numbers for each subsystem. In order to ensure orthogonality in the NSM when $M^{(p)}$ and $M^{(n)}$ are common, we start from two orthogonal Slater determinants and use the same choice of operators for both states. The second layer of EDEF splits the proton (neutron) nuclear ground state into symmetric distributions of protons (neutrons) and proton holes (neutron holes). Appendix A gives further details on the properties of these states and their corresponding coefficients, illustrated with examples.
IV Results using entropy-driven entanglement forging
@@ IV.1 Fermi-Hubbard model
In the FH model we can study systems with a wide range of entanglement and analyze the impact of the EDEF method. For this purpose, like in Sec. II.2 and Fig. 3, we simulate systems with three different central hopping values, $t_{m}=0.25t,t,2t$, and two interactions, $U=t,3t$. We obtain the ground state of these systems using two methods: the regular ADAPT-VQE and an optimized version using EDEF, as described in Sec. III. For their comparison, we focus on the infidelities of each approach with respect to the exact ground state,
$I=1-|\langle\psi_{\text{ADAPT-VQE}}|\psi_{\text{exact}}\rangle|^{2}\,.$ (15)
An alternative analysis of the ground-state energy relative error,
$\epsilon_{E}=\left|\frac{E_{\rm ADAPT-VQE}-E_{\text{exact}}}{E_{\text{exact}}}\right|,$ (16)
@@
Figure 6 compares the infidelities obtained with the regular ADAPT-VQE without entanglement forging (solid red line) and with EDEF with one layer, or “1-cut” (dashed blue line), keeping five singular values. The central hopping strength increases from left to right panels. Bottom panels correspond to a stronger interaction $U$, which causes the regular ADAPT-VQE to converge slower, but makes no difference for the 1-cut EDEF results. Horizontal lines mark the lower bound for our ansatz infidelities, given by the ones corresponding to the Schmidt decomposition of the ground state with cutoff at the fifth product state, Eq. (13). Almost all 1-cut EDEF simulations converge to this bound, highlighting the good performance of the variational optimization.
@@
Table 1: Number of qubits ($N_{q}$) per circuit used to simulate the system with regular (0 cuts) and EDEF optimized (1 or 2 cuts) ADAPT-VQE, as well as the number of iterations ($N_{\text{it}}$), relative error in the energy ($\epsilon_{E}$), infidelity ($I_{\text{conv}}$) and convergence rate ($r$) once the algorithm has either converged ($I_{\rm conv}<10^{-5}$ for ADAPT-VQE) or reached the maximum number of iterations imposed. Top: Results for FH lattices labeled as ($U$, $t_{m}$). Bottom: Results for the NSM simulations of 28Ne and 60Ti.
Figure 6 shows that for weakly linked FH lattices, $t_{m}=0.25t$, one layer of EDEF offers a clear advantage for $U=3t$ and target infidelities $\epsilon_{E}\gtrsim 10^{-4}$. For a weaker interaction $U=t$, the 1-cut EDEF infidelities also improve over regular ADAPT-VQE ones up to $I\sim 10^{-2}$ (or seven iterations). Beyond this point, ADAPT-VQE converges faster. For the canonical FH lattice, $t_{m}=t$, the entropy is still low – see Fig. 2 – and Fig. 6 indicates that 1-cut EDEF reaches almost $I\sim 10^{-2}$, although converging more slowly than the regular ADAPT-VQE. However, Fig. 6 shows that, when $t_{m}>t$, EDEF starts to fail due to the higher entropy. For these systems, reaching $I\lesssim 10^{-3}$ would require to train more than eight copies of the circuit, which would no be a clear advantage over using all qubits like in the regular ADAPT-VQE.
Table 1 (top rows) quantifies the resources and performance of the FH simulations using regular ADAPT-VQE and the EDEF optimization. It lists the number of iterations, $N_{\text{it}}$, and the infidelity, $I_{\text{conv}}$, and energy relative error, $\epsilon_{E}$, once the optimization has converged. Table 1 also presents the convergence rate, defined as
@@
Consistently with Fig. 6, the 1-cut EDEF convergence rates are higher for weaker central hopping values, but in general they are lower than for the regular ADAPT-VQE. Nonetheless, Table 1 also highlights that all 1-cut simulations require half the qubits and converge with just 14 iterations, while all regular ADAPT-VQE optimizations need all qubits and more iterations – but they reach $I\leq 10^{-5}$. Moreover, circuits with half as many qubits constrain fermion operators to be more local, and therefore involve fewer quantum gates, which is another advantage of the EDEF implementation.
@@ IV.2 Nuclear shell model
We then apply our method to study neutron-rich nuclei 28Ne and 60Ti withing the NSM. These are ideal systems for EDEF because of their low entanglement between neutrons and protons, as shown in Fig. 5. Also, they present a challenge for the regular ADAPT-VQE due to relatively large number of nucleons and single-particle states involved. In fact, in our previous work [25] we were not able to simulate 60Ti with ADAPT-VQE.
@@
We explore the performance of EDEF using various layers. In particular, we use ADAPT-VQE with circuits with $N_{q}$ (regular ADAPT-VQE), $N_{q}/2$ (EDEF with one layer), and $N_{q}/4$ (EDEF with two layers) qubits. As indicated in Fig. 4, $N_{q}$ is the number of single-particle states in the valence space, so that $N_{q}=24$ for 28Ne and $N_{q}=40$ for 60Ti. Like in Sec. IV.1, we present our results in terms of the infidelity with respect to the exact ground state obtained with each approach.
@@
The top panel of Fig. 7 presents the 28Ne infidelities as a function of the number of ADAPT-VQE iterations for the regular ADAPT-VQE (solid red line), 1-cut EDEF (dashed blue line) and 2-cut EDEF (dashed-dotted yellow line). Like for the FH model, the infidelities for one and two layers of EDEF converge very close to the lower bounds defined by the Schmidt decomposition (horizontal lines of the same color). The two EDEF variants converge faster than the regular ADAPT-VQE, meaning that with the same number of parameters, and half or one quarter of the qubits per circuit, EDEF outperforms ADAPT-VQE. The 1-cut EDEF reaches much lower infidelities, $I\sim 10^{-3}$, than the 2-cut EDEF simulation, but after significantly more iterations. Indeed, the results in Table 1 indicate that the convergence rate is similar in both cases.
@@
Figure 7: Infidelities as a function of the number of ADAPT-VQE iterations for NSM simulations of 28Ne (top) and 60Ti (bottom), using the regular ADAPT-VQE (0 cuts, solid red line) and the optimized EDEF with one (1 cut, dashed blue lines) or two (2 cuts, dashed-dotted yellow lines) layers. Horizontal lines with the same color code indicate the infidelities set by the Schmidt decompositions cut off consistently with each EDEF simulation.
@@
The bottom panel of Fig. 7 shows the results for 60Ti with one and two layers of EDEF, with the same line and color code as the upper panel. For this nucleus, standard ADAPT-VQE with no cuts exceeds our computational capabilities. While the infidelity for 1-cut EDEF shows good convergence up to $\sim 50$ iterations, at this point, when $I$ is still well above the limit set by the Schmidt decomposition, the computation becomes too slow. In contrast, the 2-cut EDEF converges to a high-quality local minimum at the 42nd iteration, with $I=0.082$ (the relative error in the energy is $\epsilon_{E}=0.16$). Consistently, Table 1 indicates a better convergence for the 2-cut simulation.
@@
In principle, EDEF can also be applied to less neutron-rich nuclei, but this implies larger proton-neutron entanglement and thus reduced efficiency. For example, 26Ne shows more complex structure in terms of product states than 28Ne. Nonetheless, the one-layer EDEF achieves only marginally poorer infidelity and relative energy error than the regular ADAPT-VQE using half as many qubits. The results for 26Ne are given in Appendix B.

### v2

@@ II.3 Nuclear shell model
In addition, nuclear entanglement can also be partly understood in terms of subshell closures and occupation numbers. Nuclei with proton or neutron orbitals that are mostly empty or fully-occupied present very low entanglement between these orbitals and the rest of states [34]. This feature motivates us to apply the EDEF approach in an iterative fashion. Firstly, we shall use a proton-neutron partition. Second, we employ an additional subpartition based on single-particle energies, separating the higher- and lower-energy states within each proton and neutron subsystem. Figure 4 illustrates the two partitions and the corresponding cuts used for 28Ne and 60Ti, which are the two isotopes of reference discussed in the following Section. For the latter nucleus, a bipartition into subsystems of the same size requires separating the energy-degenerate single-particle states of the $1p_{3/2}$ orbital.
@@ III Entropy-driven entanglement forging with ADAPT-VQE
ADAPT-VQE [59, 60, 61, 62, 63] is a variational quantum algorithm which updates iteratively a user-defined ansatz, rather than optimizing a fixed number of parameters as the UCC-VQE [24, 23, 3]. Each iteration $k$ adds a new unitary operator, $A_{k}=e^{i\theta_{k}T_{k}}$, to the ansatz, with a new parameter, $\theta_{k}$, and an Hermitian operator $T_{k}$ from a predefined operator pool. Then, all parameters, $\bm{\theta}=\{\theta_{1},\cdots,\theta_{k}\}$, are optimized simultaneously to minimize the energy,
@@
For iteration $k$, the optimization starts with the parameters values $\theta_{1},\cdots,\theta_{k-1}$ obtained in the minimization of the previous iteration, $k-1$, and with the new parameter set to zero, $\theta_{k}=0$. The operator $T_{k}$ is chosen such that the energy gradient with respect to $\theta_{k}$,
@@
is maximum. Thus, by starting from the minimum in the $(k-1)$-dimensional parameter space and maximizing the gradient in the new dimension, the algorithm gradually increases the parameter space while always decreasing the energy. ADAPT-VQE was originally proposed to solve the electronic structure of the ground state of molecules [64, 65], and since then it has been applied to a broad variety of quantum many-body systems, including lattice quantum electrodynamics [66, 67], the NSM [22, 29], algebraic nuclear models [22, 68] and nuclear pairing [69]. In addition to the novel FH simulations, we build on our previous work [29, 34] and optimize the ADAPT-VQE algorithm to solve NSM ground states with fewer quantum resources by exploiting EDEF.
@@
with the restriction that they conserve the Hamiltonian symmetries: the spin for the FH and $J_{z},T_{z}$ for the NSM. For the FH, in addition, we also include one-body operators,
$T_{s}^{r}=i(a_{r}^{{\dagger}}a_{s}-a_{s}^{{\dagger}}a_{r}),$ (12)
to the operator pool.
@@
Figure 6: Infidelity as a function of the number of ADAPT-VQE iterations for the same FH model parameters (middle hopping $t_{m}$, interaction $U$) as in Fig. 3. ADAPT-VQE without entanglement forging (solid red lines) is compared with one layer of EDEF (labelled “1 cut”, dashed blue lines). Horizontal lines indicate the infidelity minimum set by the Schmidt-decomposition cutoff.
@@
As a reference state, with the standard ADAPT-VQE we choose the lowest-energy Slater determinant in the Fock basis of the particular FH lattice or nucleus. This can be implemented with one-qubit gates under the Jordan-Wigner mapping [70]. With EDEF, however, we can exploit other physical arguments to choose a more convenient reference state, such as degeneracies in the Schmidt decomposition that strongly suggest specific symmetries. In the case of the FH model, for instance, the best choice is a symmetric distribution of up and down spins between the left ($l$) and right ($r$) partitions of the lattice,
$|\psi\rangle=$ $\lambda_{0}|l_{\uparrow\downarrow}\rangle\otimes|r_{\uparrow\downarrow}\rangle+\lambda_{1}\bigg{(}|l_{\uparrow\uparrow\downarrow}\rangle\otimes|r_{\downarrow}\rangle+|l_{\downarrow}\rangle\otimes|r_{\uparrow\uparrow\downarrow}\rangle$
$+|l_{\uparrow\downarrow\downarrow}\rangle\otimes|r_{\uparrow}\rangle+|l_{\uparrow}\rangle\otimes|r_{\uparrow\downarrow\downarrow}\rangle\bigg{)}.$ (13)
This particular choice is inspired by the Schmidt decomposition of the full state presented in Fig. 3. The analysis of the associated quantum numbers of each Schmidt vector indicates that this choice should provide an excellent reconstruction of the full state as measured by the infidelities of Fig. 2.
@@
In turn, for the NSM we use eigenstates of the $J_{z}$ operator in the proton ($p_{J_{z}}$) and neutron ($n_{J_{z}}$) sectors. The pattern for the Schmidt decomposition described in Sec. II.3 and shown in Fig. 5 suggests the following choice of reference state,
$|\psi\rangle=$ $\lambda_{0}|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle+\lambda_{1}\bigg{(}|p_{-2}\rangle\otimes|n_{2}\rangle+|p_{-1}\rangle\otimes|n_{1}\rangle$
$+|p_{0}\rangle\otimes|n_{0}\rangle+|p_{1}\rangle\otimes|n_{-1}\rangle+|p_{2}\rangle\otimes|n_{-2}\rangle\bigg{)}.$ (14)
We note that for both the FH and the NSM applications, we employ reference states that only involve two coefficients, $\lambda_{0}$ and $\lambda_{1}$, which we fix to the corresponding values of the Schmidt decomposition (see Appendix A for more details).
@@
In order to ensure orthogonality in the NSM for the states with the same ${J_{z}}^{(p)}$ and ${J_{z}}^{(n)}$, we start from two orthogonal Slater determinants and use the same choice of operators for both states. Since the operator pool preserves the quantum numbers for each subsystem, orthogonality will be maintained through the iterations. The second layer of EDEF splits the proton (neutron) nuclear ground state into symmetric distributions of protons (neutrons) and proton holes (neutron holes). Appendix A gives further details on the properties of these states and their corresponding coefficients, illustrated with the specific examples used in our optimization. Having defined the operator pools and reference states of our EDEF-accelerated ADAPT-VQE, we proceed to discuss the results of this approach for both the FH and the NSM.
IV Results
@@ IV.1 Fermi-Hubbard model
In the FH model, we can access systems with a wide range of entanglement across left and right partitions by tuning the parameter $t_{m}$. This allows us to analyze the quality of the EDEF method depending on the entanglement structure of the system. For this purpose, like in Sec. II.2 and Fig. 3, we simulate systems with three different central hopping values, $t_{m}=0.25t,t,2t$, and two interactions, $U=t,3t$. We obtain the ground state of these systems using two methods. First, we employ the full, regular ADAPT-VQE to solve for the system ground state. Second, we use an optimization based on EDEF, as described in Sec. III. To compare the efficiency of the two approaches, we focus on the infidelities of each approach with respect to the exact ground state,
$I=1-|\langle\psi_{\xi}|\psi_{\text{exact}}\rangle|^{2}\,,$ (15)
where $\xi$ can correspond to either the full ADAPT-VQE approach or to an approximate EDEF with $1$ or $2$ cuts. An alternative analysis using the ground-state energy relative error,
$\epsilon_{E}=\left|\frac{E_{\xi}-E_{\text{exact}}}{E_{\text{exact}}}\right|,$ (16)
@@
Figure 6 compares the infidelities obtained with the regular ADAPT-VQE without entanglement forging (solid red lines) and with EDEF with one layer, or “1-cut” (dashed blue lines). This approach corresponds to a reference state involving five singular values, as described in Sec. II.2. The central hopping strength increases from left to right panels. Top panels show results for $U=t$, and bottom panels correspond to a stronger interaction $U=3t$. The infidelity in logarithmic scale is shown as a function of the iteration number for both ADAPT-VQE and EDEF. In practical terms, the ADAPT-VQE simulation corresponds to a single circuit and each iteration $k$ involves the addition of a new operator $A_{k}$ and the optimization of $k$ parameters, $\theta_{1},\cdots,\theta_{k}$. We define one iteration in the 1-cut EDEF to the addition of a single new operator in either of the different subcircuits which we employ to describe the full system (see Appendix A for details on the nature and number of subcircuits). Horizontal lines mark the lower bound for our ansatz infidelities. This is given by the infidelity associated to the corresponding Schmidt decomposition of the ground state with cutoff at the fifth product state, $\chi_{\rm cut}=5$ in Eq. (13). With the degeneracy, only two different Schmidt coefficients are involved.
Even though there is no variational bound on the decrease of infidelities as there is for the energy, we find that both ADAPT-VQE and 1-cut EDEF results reduce the infidelity as the number of iterations increase. Regular ADAPT-VQE converges to $I\lesssim 1\%$ within $20$ iterations for all cases. The convergence is however relatively slower for strongly-interacting systems ($U=3t$, bottom panels) compared to less correlated systems ($U=t$, top panels). In contrast, the interaction strength $U$ makes no difference for the 1-cut EDEF results. For the more physical case of $t_{m}=t$, for instance, the 1-cut EDEF results converge to the lower bound within $14$ iterations. We stress that almost all 1-cut EDEF simulations converge to this lower bound, which highlights the good performance of the variational optimization. The exception corresponds to the extreme case of $t_{m}=2t$ and $U=3t$ (lowest right panel), which has the largest interaction values across all panels. In this case, the ansatz has a hard time converging to the exact wavefunction even for the full ADAPT-VQE simulation.
@@
Table 1: Number of qubits ($N_{q}$) per circuit used to simulate the system with regular (0 cuts) and EDEF optimized (1 or 2 cuts) ADAPT-VQE, as well as the number of iterations ($N_{\text{it}}$), relative error in the energy ($\epsilon_{E}$), infidelity ($I_{\text{conv}}$) and convergence rate ($r$) once the algorithm has either converged ($I_{\rm conv}<10^{-5}$ for ADAPT-VQE) or reached the maximum number of iterations imposed. Top: Results for FH lattices labeled as ($t_{m}$, $U$). Bottom: Results for the NSM simulations of 28Ne and 60Ti.
The left panels of Fig. 6 show that for weakly linked FH lattices, $t_{m}=0.25t$, one layer of EDEF offers a clear advantage for $U=3t$ and target infidelities $\epsilon_{E}\gtrsim 10^{-4}$. For a weaker interaction $U=t$, the 1-cut EDEF infidelities also improve over regular ADAPT-VQE up to $I\sim 10^{-2}$ (or seven iterations). Beyond this point, ADAPT-VQE converges faster. For the canonical FH lattice, $t_{m}=t$ (central panels), the entropy is still low – see Fig. 2 – and Fig. 6 indicates that 1-cut EDEF reaches almost $I\sim 10^{-2}$, although converging more slowly than the regular ADAPT-VQE. Figure 6 also shows that, when $t_{m}>t$ (right panels), EDEF starts to underperform substantially compared to ADAPT-VQE. This can be associated to a larger entanglement entropy. For these systems, reaching $I\lesssim 10^{-3}$ may require to train more than eight copies of the circuit which, at this system size, does not provide a clear performance advantage over using all the qubits like in the regular ADAPT-VQE – although the local optimization could still provide some benefits. We postpone a thorough analysis of these issues for further work.
Table 1 (top rows) quantifies the resources and performance of the FH simulations using regular ADAPT-VQE and the EDEF optimization. It lists the number of iterations, $N_{\text{it}}$, the infidelity, $I_{\text{conv}}$, and energy relative error, $\epsilon_{E}$, once the optimization has converged. As a convergence criterion, we stop our ADAPT-VQE simulations when the infidelity falls below $I<10^{-5}$. For EDEF, convergence is often found above this threshold value, when gradients reach very small values and updates do not improve the energy any further. In other cases, particularly for the more demanding NSM simulations, we stop the simulations arbitrarily after $100$ iterations.
Table 1 presents a proxy for the state of the convergence procedure, the so-called convergence rate, which we define as
@@
Consistently with Fig. 6, the 1-cut EDEF convergence rates are higher for weaker central hopping values, but in general they are lower than for the regular ADAPT-VQE. Nonetheless, Table 1 also highlights that all 1-cut simulations require $4$ qubits and converge with just $14$ iterations, while all regular ADAPT-VQE optimizations need all $8$ qubits and $16$ or more iterations. In contrast, the infidelity of the converged results reaches a high quality for ADAPT-VQE, with $I\leq 10^{-5}$ across different values of $t_{m}$ and $U$. The best EDEF simulations reach at most $I\approx 10^{-4}$, and often reach worse values. Having said that, circuits with half as many qubits also constrain fermion operators to be more local. Therefore, in addition to reducing the amount of qubits needed, these require fewer quantum gates, which is another advantage of the EDEF implementation, as we analyse in Sec. IV.3.
@@ IV.2 Nuclear shell model
We now apply the EDEF method to study neutron-rich isotopes 28Ne and 60Ti within the NSM. These are ideal systems for EDEF, because of the low entanglement between the neutron and proton partitions, as shown in Fig. 5. Moreover, these are challenging isotopes from the regular ADAPT-VQE perspective, because they involve a relatively large number of nucleons and single-particle states. In fact, in our previous work [29] we were not able to simulate 60Ti with ADAPT-VQE because of the large memory required to encode the corresponding state vectors.
@@
We explore the performance of EDEF using various layers. In particular, we use ADAPT-VQE with circuits with $N_{q}$ (regular ADAPT-VQE), $N_{q}/2$ (EDEF with one cut), and $N_{q}/4$ (EDEF with two cuts) qubits. $N_{q}$ is the number of single-particle states in the valence space, which in the mapping described in Fig. 4 also corresponds to the number of qubits. In the $sd$ shell, we require $N_{q}=24$ qubits to simulate 28Ne, whereas for 60Ti in the $pf$ shell, $N_{q}=40$ qubits are needed. Similarly to Sec. IV.1, we quantify the quality of both the full ADAPT-VQE and the EDEF approach in terms of the infidelity with respect to the exact ground state obtained with each approach.
@@
The top panel of Fig. 7 presents the evolution of the infidelity of the ground-state wave function of 28Ne as a function of the number of ADAPT-VQE iterations for the regular ADAPT-VQE (solid red line), 1-cut EDEF (dashed blue line) and 2-cut EDEF (dashed-dotted yellow line). As for the FH model, the infidelities for one and two layers of EDEF converge very close to the lower bounds defined by the Schmidt decomposition (horizontal lines of the corresponding color). The two EDEF variants converge faster than the regular ADAPT-VQE, meaning that with the same number of parameters, and half or one quarter of the qubits per circuit, EDEF outperforms ADAPT-VQE. We note that the full ADAPT-VQE simulations increase the infidelity with respect to the exact ground state between iterations $15$ and $60$, where a sudden drop occurs and the decline of infidelity resumes. This is possible because ADAPT-VQE is a variational optimisation process based on the minimization of energies, so the infidelity can increase when not too close to the target energy – unless the ground state space is degenerate or the energy spectrum is gapless, the infidelity must eventually decrease as the error in energy approaches zero. The 1-cut EDEF shows a much milder increase of infidelity between iterations $15$ and $25$. It also converges to much lower infidelities, $I\approx 10^{-3}$, than the overall 2-cut EDEF simulation. This happens, however, after a significantly larger number of iterations. Indeed, the results in Table 1 indicate that the convergence rate is similar in both cases.
@@
Figure 7: Infidelities as a function of the number of ADAPT-VQE iterations for NSM simulations of 28Ne (top panel) and 60Ti (bottom panel), using the regular ADAPT-VQE (0 cuts, solid red line) and the optimized EDEF with one (1 cut, dashed blue lines) or two (2 cuts, dashed-dotted yellow lines) cuts. Horizontal lines with the same colour code indicate the infidelities set by the Schmidt decompositions cut off consistently with each EDEF simulation.
@@
The bottom panel of Fig. 7 shows the results for 60Ti with one and two layers of EDEF, with the same line and color code as the top panel. For this nucleus, standard ADAPT-VQE with no cuts exceeds our computational capabilities and is thus not shown here. The infidelity for 1-cut EDEF shows good convergence up to $\sim 50$ iterations. At this point, when $I$ is still well above the limit set by the Schmidt decomposition, the computation becomes too slow for our current computational resources and we stop it. In contrast, the 2-cut EDEF requires significantly less resources. We find that this 2-cut EDEF simulation is able to converge to a high-quality local minimum at the 42nd iteration, with an infidelity of $I=0.082$. The corresponding relative error in the energy is $\epsilon_{E}=0.16$. Consistently, Table 1 indicates a better convergence in terms of infidelities, energies and rates for the 2-cut simulation as opposed to the 1-cut approach for this isotope.
@@
In principle, nothing precludes the application of EDEF to less neutron-rich nuclei. However, in general we expect these to have larger proton-neutron entanglement and thus the EDEF approach to work less efficiently. For instance, Fig. 5 shows that 26Ne has a more complex structure in terms of product states than 28Ne, with significantly larger $i=7$ and $8$ Schmidt coefficients. Nonetheless, the one-layer EDEF achieves only marginally poorer infidelity and relative energy error than the regular ADAPT-VQE, but using half as many qubits and fewer resources. We provide the ADAPT-VQE and EDEF results for 26Ne in Appendix C.